Which ordered pair is the solution of the system?

6x - 6y = 2
3x + 9y = -7

(1/3, 2/3)
(2/3, -1/3)
(-1/3, -2/3)
(-2/3, 1/3)

Answer by stanbon(75887)

Which ordered pair is the solution of the system?
6x - 6y = 2
3x + 9y = -7
----
Modify::
6x - 6y = 2
6x + 18y = -14
------
24y = -16
y = -2/3
-----
Solve for "x"::
3x + 9(-2/3) = -7
3x - 18/3 = -21/3
---
3x = -1
-----
x = -1/3
---------------------
Cheers,
Stan H.
----------------
(1/3, 2/3)
(2/3, -1/3)
(-1/3, -2/3)
(-2/3, 1/3)